OpenStudy (anonymous):

Evaluate the following expression using the values given: Find -a2 - 3b3 + c2 + 2b3 - c2 if a = 3, b = 2, and c = -3. -67 -17 -31 1

OpenStudy (unklerhaukus):

\[-a^2 - 3b^3 + c^2 + 2b^3 - c^2%\\=-(3)^2 - 3(2)^3 + (-3)^2 + 2(2)^3 - (-3)^2 \] simplify the equation before substituting in the values

OpenStudy (unklerhaukus):

\[-a^2 - 3b^3 + c^2 + 2b^3 - c^2\\=-a^2 + 2b^3- 3b^3 + c^2 - c^2\\ =-a^2+(2-3)b^3+(1-1)c^2\\ =\]